OpenStudy (anonymous):

Solve for x: log4x+log4(x+1)=7

OpenStudy (anonymous):

this is \[\log_4(x)+\log_4(x+1)=7\] right? then \[\log_4(x(x+1))=7\] so \[x^2+x=7^4\] yikes

OpenStudy (anonymous):

typo should be \[x^2+x=4^7\] double yikes

OpenStudy (anonymous):

if this is really the problem, your job is to solve \[x^2+x-16384=0\] using the quadratic formula. glad it is you and not me

OpenStudy (anonymous):

i don't suppose it might be \[\log_4(x)+\log_4(x+1)=\log_4(7)\] is it?

OpenStudy (anonymous):

actually thats what it is i forgot to put exponents thanks for correcting me